# ResQ: Mixed-Precision Quantization of Large Language Models with Low-Rank Residuals

### v1

@@ Acknowledgment
Appendix A Appendix
A.1 Proof of Theorem 7
@@
Proof of lemma A.1 can be found in Li et al. (2024). From this lemma we obtain that the quantization error $\lVert{\bm{R}}-Q({\bm{R}})\rVert_{F}$ is bounded by the magnitude of the tensor quantized $\lVert{\bm{R}}\rVert_{F}$. Now for our use case of mixed precision quantization where the low-precision component is quantized to $L$ bits and high precision component is quantized to $H$ bits, we write the quantization error again below,
@@
$\begin{split}\lVert{{\bm{X}}}-{{\bm{X}}_{q}}\rVert_{F}&\leq\frac{\sqrt{\text{log(size}({\bm{X}}{\bm{U}}_{l}))\pi}}{2^{L-1}-1}\mathbb{E}||{\bm{X}}{\bm{U}}_{l}||_{F}\\ &+\frac{\sqrt{\text{log(size}({\bm{X}}{\bm{U}}_{h}))\pi}}{2^{H-1}-1}\mathbb{E}||{\bm{X}}{\bm{U}}_{h}||_{F}\\ &=\frac{\sqrt{\text{log(size}({\bm{X}}{\bm{P}}_{l}))\pi}}{2^{L-1}-1}\mathbb{E}||{\bm{X}}{\bm{P}}_{l}||_{F}\\ &+\frac{\sqrt{\text{log(size}({\bm{X}}{\bm{P}}_{h}))\pi}}{2^{H-1}-1}\mathbb{E}||{\bm{X}}{\bm{P}}_{h}||_{F}\\ &=\frac{\sqrt{\text{log(size}({\bm{X}}{\bm{P}}_{l}))\pi}}{2^{L-1}-1}\mathbb{E}||\text{tr}({\bm{X}}{\bm{P}}_{l}{\bm{P}}^{\top}_{l}{\bm{X}}^{\top})||_{F}\\ &+\frac{\sqrt{\text{log(size}({\bm{X}}{\bm{P}}_{h})\pi}}{2^{H-1}-1}\mathbb{E}||\text{tr}({\bm{X}}{\bm{P}}_{h}{\bm{P}}^{\top}_{h}{\bm{X}}^{\top})||_{F}\end{split}$ (11)
Model Method Llama-2 family
0-shot common sense reasoning tasks 0-shot MMLU tasks
ARC-c ARC-e BoolQ HellaS OBQA PIQA SIQA WinoG Avg. humanities Other SocialS STEM Avg.
($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$)
Llama-2-7b 16-bit 46.3 74.6 77.8 75.9 44.2 79.2 46.1 69.1 64.1 38.9 45.9 46.0 33.4 41.1
[…]
GPTQ 24.8 31.4 55.4 30.6 25.6 55.8 34.2 53.3 38.9 24.7 24.5 22.7 23.2 23.8
SmoothQuant+ 29.3 47.1 56.8 48.6 31.8 65.5 37.2 52.4 46.1 25.0 24.5 24.1 23.4 24.2
QUIK 39.8 63.7 68.9 68.3 37.8 72.9 42.1 62.4 57.0 26.9 29.6 28.8 25.8 27.8
QuaRot 41.5 71.4 73.2 73.2 40.6 76.9 43.6 65.6 60.7 31.2 35.1 34.6 28.2 32.3
SpinQuant 43.6 71.3 73.8 73.2 40.4 76.0 44.1 65.4 61.0 33.9 38.5 37.5 29.5 34.8
ResQ 44.0 72.6 75.3 74.0 41.0 77.9 43.9 66.9 62.0 35.9 40.9 42.2 32.2 37.7
Llama-2-13b 16-bit 49.1 77.4 80.5 79.4 45.2 80.7 47.2 72.1 66.5 47.9 59.3 61.0 42.4 52.7
[…]
GPTQ 23.6 31.1 38.7 27.2 26.8 53.6 35.8 49.8 33.8 25.0 25.4 23.7 25.1 24.8
SmoothQuant+ 34.5 55.6 62.9 62.5 32.4 70.1 38.7 55.6 51.0 25.7 26.1 27.3 27.3 26.6
QUIK 43.7 68.0 71.3 73.3 40.0 75.7 45.1 64.6 60.2 34.7 40.6 39.8 31.8 36.7
QuaRot 46.9 74.9 76.6 75.8 42.6 79.1 45.5 69.0 63.8 43.8 53.6 54.0 39.4 47.7
SpinQuant 49.0 76.3 78.2 77.1 42.8 79.3 46.3 69.5 64.8 43.5 53.1 55.4 39.1 47.8
ResQ 49.1 76.1 79.7 77.9 43.6 79.1 46.6 69.9 65.2 45.3 56.0 58.0 41.0 50.1
Model Method Llama-3 family
0-shot common sense reasoning tasks 0-shot MMLU tasks
ARC-c ARC-e BoolQ HellaS OBQA PIQA SIQA WinoG Avg. humanities Other SocialS STEM Avg.
($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$)
Llama-3-8b 16-bit 53.2 77.1 81.1 79.2 44.8 80.9 47.0 73.4 67.1 55.0 70.6 73.2 53.7 63.1
[…]
GPTQ 24.7 37.7 44.3 36.8 27.0 57.6 36.4 53.8 39.8 24.7 23.9 22.8 21.8 23.3
SmoothQuant+ 27.5 42.0 50.7 44.9 28.8 59.0 35.9 50.9 42.5 25.4 25.5 24.5 23.4 24.7
QUIK 33.6 56.4 60.5 61.5 33.2 68.7 39.9 59.0 51.6 30.0 34.0 34.8 32.1 32.7
QuaRot 45.1 70.4 73.8 74.7 42.6 76.6 45.1 68.5 62.1 47.8 59.1 61.4 44.3 53.2
SpinQuant 48.0 75.4 75.8 75.4 43.8 77.5 45.0 69.2 63.8 49.8 63.3 65.0 46.8 56.2
ResQ 49.2 75.0 72.5 76.5 43.0 78.3 45.8 71.0 63.9 50.6 64.4 65.8 48.1 57.2
Llama-3-70b 16-bit 64.2 85.9 85.3 84.9 48.6 84.4 50.8 80.6 73.1 67.6 81.5 86.8 68.4 76.1
[…]
GPTQ 25.9 26.0 37.9 26.2 28.6 50.4 34.3 49.9 34.9 27.1 24.3 24.0 26.5 25.5
SmoothQuant+ - - - - - - - - - - - - - -
QUIK 44.5 68.9 60.7 75.0 36.4 76.1 43.2 60.4 58.2 46.6 56.4 58.0 43.6 51.1
QuaRot 53.7 74.5 81.6 81.1 46.6 81.0 46.8 75.2 67.6 55.7 72.5 75.8 57.3 65.3
SpinQuant 52.0 77.3 81.7 75.6 43.8 78.8 43.4 72.8 65.7 50.7 67.0 68.1 51.9 59.4
ResQ 61.4 84.3 83.9 83.5 46.0 83.1 48.6 78.3 71.1 64.9 79.9 84.9 66.1 74.0
Model Method Llama-3.2 family
0-shot common sense reasoning tasks 0-shot MMLU tasks
ARC-c ARC-e BoolQ HellaS OBQA PIQA SIQA WinoG Avg. humanities Other SocialS STEM Avg.
($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$) ($\uparrow$)
Llama-3.2-1b 16-bit 36.5 60.6 63.4 63.6 37.4 74.5 42.8 60.1 54.9 34.8 41.1 39.9 32.0 36.9
[…]
GPTQ 24.7 32.7 52.3 30.7 23.6 54.3 34.4 51.1 38.0 24.7 25.1 25.5 24.5 24.9
SmoothQuant+ 23.3 30.1 52.9 31.3 26.6 54.2 34.5 51.2 38.0 23.9 24.1 25.0 23.5 24.1
QUIK 27.4 46.0 55.0 46.0 26.4 62.4 38.6 52.6 44.3 25.6 25.6 24.6 24.5 25.1
QuaRot 30.0 51.4 59.1 54.0 34.2 66.7 39.6 57.1 49.0 25.4 26.9 25.4 24.4 25.5
SpinQuant 32.3 51.8 59.3 55.4 30.4 67.7 38.6 54.7 48.8 25.4 27.6 24.2 25.3 25.6
ResQ 34.0 54.2 57.0 57.3 31.2 69.4 41.0 56.8 50.1 28.3 30.5 31.3 27.6 29.4
Llama-3.2-3b 16-bit 46.2 71.7 73.1 73.7 43.4 77.4 47.2 69.1 62.7 48.9 62.9 62.3 45.2 54.8
[…]
GPTQ 27.0 27.0 48.8 44.4 27.8 59.1 37.1 51.5 40.3 24.9 24.5 25.7 24.0 24.8
SmoothQuant+ 25.3 33.1 47.8 37.7 25.2 56.2 35.8 50.9 39.0 25.4 26.6 26.4 25.3 25.9
QUIK 32.9 50.1 52.6 59.1 33.2 68.7 40.3 53.0 48.8 29.0 33.2 31.9 30.3 31.1
QuaRot 38.6 59.0 65.9 66.5 35.8 74.4 43.1 65.2 56.1 38.5 47.3 46.7 35.3 42.0
SpinQuant 38.9 64.8 68.0 69.1 39.4 74.9 45.1 62.9 57.9 37.0 49.4 50.5 39.9 44.2
ResQ 43.1 65.6 68.8 70.5 38.4 75.1 45.6 64.8 59.0 44.7 57.0 56.5 41.0 49.8
Table 5: Accuracy on eight 0-shot common sense reasoning tasks including ARC-challenge, ARC-easy, BoolQ, HellaSwag, Openbook QA, PIQA, SIQA, and WinoGrande and 0-shot massive multitask language understanding tasks across four subjects: STEM, Humanities, Social Sciences, and MMLU-other, for the Llama-2, Llama-3 and Llama-3.2 families when quantized to W/A/KV = 4/4/4. Results of all techniques were obtained using their official codebase. Our work ResQ and QUIK Ashkboos et al. (2023) keep $\nicefrac{{1}}{{8}}$th of channels in 8-bit. All techniques except RTN use GPTQ Frantar et al. (2022).
@@
$\begin{split}\lVert{{\bm{X}}}-{{\bm{X}}_{q}}\rVert_{F}&\leq\frac{\sqrt{\text{log(d-r)}\pi}}{2^{L-1}-1}(\mathbb{E}\lVert{{\bm{X}}}\rVert_{F}-\mathbb{E}\lVert{{\bm{X}}{\bm{P}}_{h}}\rVert_{F})\\ &+\frac{\sqrt{\text{log(r)}\pi}}{2^{H-1}-1}\mathbb{E}\lVert{{\bm{X}}{\bm{P}}_{h}}\rVert_{F}\\ &=\frac{\sqrt{\text{log(d-r)}\pi}}{2^{L-1}-1}\mathbb{E}\lVert{{\bm{X}}}\rVert_{F}\\ &-(\frac{\sqrt{\text{log(d-r)}\pi}}{2^{L-1}-1}-\frac{\sqrt{\text{log(r)}\pi}}{2^{H-1}-1})\mathbb{E}\lVert{{\bm{X}}{\bm{P}}_{h}}\rVert_{F}\end{split}$ (12)
@@
Figure 6: Input activation distributions of attention and FFN layers, for baseline (a and c) and ResQ (a and b).
@@
Llama-3-8b 17m 35m 1h41m
Llama-2-7b 15m 33m 1h37m
Llama-2-13b 23m 1h 3h42m
A.2 Distribution of activations
The distribution of activations after projection by ${\bm{U}}$ is shown in Figure 6. The formulation of ${\bm{U}}$ ensures that the final $r$ channels in the activation map comprise of coefficients along basis vectors with maximum activation variance. Consequently, keep those channels in high precision minimizes quantization error. The remaining channels are more amenable to quantization due to the application of random rotations which suppress outlier values.
A.3 Complete results of main result tables
Task by task results of the Table 1 in main paper is shown in Table 5. As expected, ResQ achieves superior performance than baselines across the series of common sense reasoning and MMLU tasks.
A.4 Additional implementation details
@@
In this work, obtaining the projection matrices and quantization of weights for all the models is performed on a single NVIDIA A100 80GB GPUs. Time taken by ResQ compared with other approaches is shown in Table 6. Evaluation on various benchmarks for all the models is also done on a single NVIDIA A100 GPU with the sole exception of Llama-3-70b which requires 4 GPUs for evaluation. We use lm_evaluation_harness version 0.4.5 Gao et al. (2024) and LongBench Bai et al. (2024) for all the evaluation tasks. For Arc-c/e, Hellaswag, OpenBook QA, PIQA tasks we report acc_norm while for BoolQ, SIQA and Winogrande we report acc.
@@
For calibration data, we use 512 randomly choses samples for Wikitext to obtain the projection matrices. While for GPTQ we use 128 randomly choses samples from Wiktiext following the original work Frantar et al. (2022).

### v2

@@ Acknowledgements
Appendix A Proof of Theorem 7
@@
Proof of lemma A.1 can be found in (Li et al., 2024). From this lemma we obtain that the quantization error $\lVert{\bm{R}}-Q({\bm{R}})\rVert_{F}$ is bounded by the magnitude of the tensor quantized $\lVert{\bm{R}}\rVert_{F}$. Now for our use case of mixed precision quantization where the low-precision component is quantized to $L$ bits and high precision component is quantized to $H$ bits, we write the quantization error again below,
@@
$\small\begin{split}\lVert{{\bm{X}}}-{{\bm{X}}_{q}}\rVert_{F}&\leq\frac{\sqrt{\text{log(size}({\bm{X}}{\bm{U}}_{l}))\pi}}{2^{L-1}-1}\mathbb{E}||{\bm{X}}{\bm{U}}_{l}||_{F}\\ &+\frac{\sqrt{\text{log(size}({\bm{X}}{\bm{U}}_{h}))\pi}}{2^{H-1}-1}\mathbb{E}||{\bm{X}}{\bm{U}}_{h}||_{F}\\ &=\frac{\sqrt{\text{log(size}({\bm{X}}{\bm{P}}_{l}))\pi}}{2^{L-1}-1}\mathbb{E}||{\bm{X}}{\bm{P}}_{l}||_{F}\\ &+\frac{\sqrt{\text{log(size}({\bm{X}}{\bm{P}}_{h}))\pi}}{2^{H-1}-1}\mathbb{E}||{\bm{X}}{\bm{P}}_{h}||_{F}\\ &=\frac{\sqrt{\text{log(size}({\bm{X}}{\bm{P}}_{l}))\pi}}{2^{L-1}-1}\mathbb{E}||\text{tr}({\bm{X}}{\bm{P}}_{l}{\bm{P}}^{\top}_{l}{\bm{X}}^{\top})||_{F}\\ &+\frac{\sqrt{\text{log(size}({\bm{X}}{\bm{P}}_{h})\pi}}{2^{H-1}-1}\mathbb{E}||\text{tr}({\bm{X}}{\bm{P}}_{h}{\bm{P}}^{\top}_{h}{\bm{X}}^{\top})||_{F}\end{split}$ (11)
@@
$\small\begin{split}\lVert{{\bm{X}}}-{{\bm{X}}_{q}}\rVert_{F}&\leq\frac{\sqrt{\text{log(d-r)}\pi}}{2^{L-1}-1}(\mathbb{E}\lVert{{\bm{X}}}\rVert_{F}-\mathbb{E}\lVert{{\bm{X}}{\bm{P}}_{h}}\rVert_{F})\\ &+\frac{\sqrt{\text{log(r)}\pi}}{2^{H-1}-1}\mathbb{E}\lVert{{\bm{X}}{\bm{P}}_{h}}\rVert_{F}\\ &=\frac{\sqrt{\text{log(d-r)}\pi}}{2^{L-1}-1}\mathbb{E}\lVert{{\bm{X}}}\rVert_{F}\\ &-(\frac{\sqrt{\text{log(d-r)}\pi}}{2^{L-1}-1}-\frac{\sqrt{\text{log(r)}\pi}}{2^{H-1}-1})\mathbb{E}\lVert{{\bm{X}}{\bm{P}}_{h}}\rVert_{F}\end{split}$ (12)
@@
Figure 7: Input activation distributions of attention and FFN layers, for baseline (a and c) and ResQ (b and d).
Appendix B Distribution of activations
The distribution of activations after projection by ${\bm{U}}$ is shown in Figure 7. The formulation of ${\bm{U}}$ ensures that the final $r$ channels in the activation map comprise of coefficients along bases with maximum activation variance. Consequently, keep those channels in high precision minimizes quantization error. The remaining channels are more amenable to quantization due to the application of random rotations which suppress outlier values.
@@ Appendix B Distribution of activations
Meta-Llama-3-8B 17m 35m 1h41m
Llama-2-7b-hf 15m 33m 1h37m
Llama-2-13b-hf 23m 1h 3h42m
Appendix C Additional implementation details
@@
In this work, obtaining the projection matrices and quantization of weights for all the models is performed on a single NVIDIA A100 80GB GPUs. Time taken by ResQ compared with other approaches is shown in Table 6. Evaluation on various benchmarks for all the models is also done on a single NVIDIA A100 GPU with the sole exception of Meta-Llama-3-70b which requires 4 GPUs for evaluation. We use lm_evaluation_harness version 0.4.5 (Gao et al., 2024) and LongBench (Bai et al., 2024) for all the evaluation tasks. For Arc-c/e, Hellaswag, OpenBook QA, PIQA tasks we report acc_norm while for BoolQ, SIQA and Winogrande we report acc.
@@
For calibration data, we use 512 randomly choses samples for Wikitext to obtain the projection matrices. While for GPTQ we use 128 randomly choses samples from Wiktiext following the original work (Frantar et al., 2022).